# Error correction of parity-encoding-based annealing through post-readout decoding

### 13 Feb 2024 v1

@@ I Introduction
In the ME scheme, each fully connected spin (logical spin) is encoded on an associated chain of physical spins, and physical spins belonging to the same chain are connected by strong ferromagnetic interactions to align them energetically. Long-range interactions between logical spins are replaced by short-range interactions between chains of physical spins. As a result, a general Ising problem that is defined on a fully connected graph of many nodes can be described by an equivalent problem that is defined on a sparse graph with a larger number of nodes Vinci et al. (2015). In contrast, in the PE scheme, which was first proposed by Sourlas Sourlas (2005) and later by Lechner, Hauke, and Zoller (LHZ) Lechner et al. (2015), the parity of each pair of logical spins is encoded on an associated physical spin that is arranged on a two-dimensional lattice. Four nearest-neighbor physical spins are coupled by four-body interactions and constitute a PE spin network. Although this approach can avoid long-range interactions, it requires four-body interactions, which appears physically unnatural and unfeasible. However, Puri et al. later suggested that the PE spin network is realizable by mapping it to a network of two-photon-driven Kerr parametric oscillators (KPOs) with only local interactions, where a four-body interaction is mapped to a four-wave mixing of four-mode oscillator fields in the Josephson junction coupler Puri et al. (2017). Such a network of KPOs can be implemented using a realistic-circuit QED platform employing superconducting devices, which has been studied extensively to realize quantum computing with the current noisy intermediate-scale quantum devices Albash et al. (2016); Lechner et al. (2015). Chancellor, Zohren, and Warburton independently designed a four-body coupling of a superconducting flux quantum bit (qubit), which is a more naive model of the quantum spin, using additional ancillary qubits Chancellor et al. (2017). In the PE scheme, the optimization problem is encoded in the dataset of the magnitudes as well as orientations of the local fields that act independently on each physical spin, whereas problem-independent four-body couplings among the neighboring physical spins enforce the parity constraints.
@@
In this study, we investigated another aspect of the PE scheme, that is, its error-correcting capability, by comparing it with that of the quantum annealing correction (QAC) scheme Jordan et al. (2006); Young et al. (2013a); Pudenz et al. (2014, 2015); Vinci et al. (2015); Bookatz et al. (2015); Matsuura et al. (2016, 2017); Pearson et al. (2019). QAC is a QA system that is merged with the classical repetition code to mitigate the errors arising in the QA algorithm. Its error-correcting capability has been discussed in terms of that of the classical repetition code. The error-correcting capability of the PE scheme is more subtle. LHZ suggested that the PE scheme includes intrinsic fault tolerance because the logical spins are encoded redundantly and nonlocally in the physical spins. Pastawski and Preskill suggested that the PE scheme can be viewed as a classical low-density parity-check (LDPC) code that makes the scheme highly robust against weakly correlated spin-flip noise Pastawski and Preskill (2016). However, a numerical study using simulated QA and parallel tempering showed no clear evidence that intrinsic fault tolerance can be used to improve the sampling efficiency of the ground state of logical spins Albash et al. (2016). To elucidate this inconsistency, we conducted discrete-time Markov chain Monte Carlo (DT-MCMC) simulations to sample the sequence of state of physical spins under various constant simulation parameters. As our MCMC is reversible, the sequence converges with the stationary distribution related to the Boltzmann distribution. Through the MCMC simulation, we stored the states of the physical spins sampled during the MCMC iteration in the buffer memory. After completing the simulation, we decoded the physical state using one-step majority vote decoding (one-step MVD) Massey (1963, 1968); Laferriere (1977); Clerk and Cain (1981); Shu and Costello (2004) to infer the state of the logical spins and calculated their energies for all the elements in the stored sequence of the physical states. Their energy spectra indicated that a substantial number of non-code states contributed to the logical ground state after the decoding. In particular, we found that the optimal simulation parameters for sampling the physical ground state using no post-readout decoding and the logical ground states using one-step MVD are very different.
@@ II.1 Soft annealing
$\stackrel{{\scriptstyle[}}{{j}}=1]{N}{\bigoplus}H_{ij}c_{j}=0\;\left(i=1,\ldots,N-K\right)$ (2)
@@
This defines the Boolean matrix $\boldsymbol{M}$ called the “parity constraint matrix” that is associated with the parity check matrix $\boldsymbol{H}$ of the code. $M_{k_{1}\cdots k_{q}}^{l}$ is $1$ if each parity equation consists of the Boolean sum of $c_{k_{l}}\left(l=1,\ldots,p\right)$. The physical state $\boldsymbol{z}$ that satisfies Eq. (4) is called a code state. For the general physical state $\boldsymbol{z}^{\prime}$, we refer to $S_{i}\left(\boldsymbol{z}^{\prime}\right)$ syndrome, which is the classical counterpart of the stabilizer operator defined in terms of the quantum correcting codes Gottesman (1996); Jordan et al. (2006). In some studies, it is also called the parity check in the context of the communication application Massey (1963, 1968).
@@
$P\left(\boldsymbol{z}|\boldsymbol{J}\right)$ $=$ $\frac{P\left(\boldsymbol{J}|\boldsymbol{z}\right)P\left(\boldsymbol{z}\right)P\left(\boldsymbol{Z}\right)}{\sum_{\boldsymbol{\sigma}}P\left(\boldsymbol{J}|\boldsymbol{z}\right)P\left(\boldsymbol{z}\right)P\left(\boldsymbol{Z}\right)}$ (5)
@@
where $\kappa$ is a constant independent of $\boldsymbol{z}$, which is to be determined by the normalization condition $\mathop{\sum_{\boldsymbol{\gamma}}P\left(\boldsymbol{z}|\boldsymbol{J}\right)=1}$. $P\left(\boldsymbol{z}\right)$ and $P\left(\boldsymbol{Z}\right)$ are the prior probabilities for the physical spin state $\boldsymbol{z}$ and logical spin state $\boldsymbol{Z}$, respectively Vicente et al. (2003). In the present scenario, we assume an informative prior for the parity constraints of the physical spin state $\boldsymbol{z}$ but ignorance prior for the logical spin state $\boldsymbol{Z}$, that is,
@@ II.4 QAC
which commutes with $\hat{H}^{enc}\left(\hat{\boldsymbol{z}}\right)$ and is introduced to take advantage of the prior knowledge of $K-1$ sets of the parity constraints on the $K$ physical spin variables $\hat{z}_{i\,k}$ for $k=1,\ldots,K$ in the Bayes inference. $\hat{H}^{pen}\left(\hat{\boldsymbol{z}}\right)$ enforces the alignment of the $K$ physical spins in each logical group. The Hamiltonian $\hat{H}^{two}\left(\hat{\boldsymbol{z}}\right)$ introduces correlations between the physical spins in different logical groups, which are absent in the Hamiltonian $\hat{H}^{phys}\left(\hat{\boldsymbol{z}}\right)$ in the communication scenario. In the limit $\gamma\rightarrow 0$, $\hat{H}^{QAC}\left(\hat{\boldsymbol{z}}\right)$ approaches the Hamiltonian for $K$ independent copies of the $N$ spin system, each of which is described by the logical Hamiltonian $\left.\hat{H}^{logi}\left(\hat{\boldsymbol{Z}}\right)\right|_{\hat{Z}_{i}\mapsto\hat{z}_{i,k}}$. Hence, we refer to a set of $N$ physical spins specified by $k$ as a replica. In contrast, in the limit $\beta\rightarrow 0$, $\hat{H}^{QAC}\left(\hat{\boldsymbol{z}}\right)$ approaches the Hamiltonian $\hat{H}^{pen}\left(\hat{\boldsymbol{z}}\right)$ that energetically penalizes the non-code state of the $\left[N,1\right]$ repetition spin-flip code. In this case, it should be noted that if we select the weight-two stabilizer in Eq. (38), $\hat{H}^{QAC}\left(\hat{\boldsymbol{z}}\right)$ is formally identical to that in simulated QA (SQA) using the path-integral Monte Carlo method Marto?ak et al. (2002); Heim et al. (2015); Waidyasooriya and Hariyama (2020); Hu and Wang (2021). This implies that SQA may be re-investigated from the viewpoint of the classical error-correcting code. We can describe the quantum dynamics of QAC by combining the standard transverse-field driver Hamiltonian based on local Pauli-$x$ operators
@@ III Decoding of post-readout spin state
$\boldsymbol{Z}^{*}$ $=$ $\underset{\boldsymbol{Z}\in\left\{\pm 1\right\}^{K}}{\arg\max}P\left(\boldsymbol{Z}|\boldsymbol{J}\right)=\underset{\boldsymbol{Z}\in\left\{\pm 1\right\}^{K}}{\arg\max}L\left(\boldsymbol{Z},\boldsymbol{J}\right)$ (47)
$=$ $\underset{\boldsymbol{Z}\in\left\{\pm 1\right\}^{K}}{\arg\min}H^{logi}\left(\boldsymbol{Z}\right)=\boldsymbol{Z}_{g}$
@@
$\boldsymbol{z}^{*}$ $=$ $\underset{\boldsymbol{z}\in\left\{\pm 1\right\}^{N}}{\arg\max}P\left(\boldsymbol{z}|\boldsymbol{J}\right)=\underset{\boldsymbol{z}\in\left\{\pm 1\right\}^{N}}{\arg\max}L\left(\boldsymbol{z},\boldsymbol{J}\right)$ (48)
$=$ $\underset{\boldsymbol{z}\in\left\{\pm 1\right\}^{N}}{\arg\min}H^{phys}\left(\boldsymbol{z}\right)=\boldsymbol{z}_{g}$
@@
In the following, we focus on the case where the target states are the logical ground states $\boldsymbol{Z}_{g}$ and associated physical code state $\boldsymbol{z}_{g}=\boldsymbol{Z}_{g}\otimes\boldsymbol{Z}_{g}$. This is because our main objective is the development of an efficient solver for combinatorial optimization problems by taking advantage of nature-inspired approaches. It is assumed that we have a method that can statistically sample the neighboring states of the physical target state $\boldsymbol{r}\sim\boldsymbol{z}_{g}$, such as simulated annealing based on Monte Carlo simulation and QA. The readout state $\boldsymbol{r}$ may be a non-code state. We address the decoding of the interim readout $\boldsymbol{r}$ to infer the target state $\boldsymbol{z}^{*}$ by taking advantage of the informative prior for parity constraints incorporated by the error-correcting codes. Let us consider the following scenario. Given $\boldsymbol{J}$ and $\boldsymbol{h}$, we statistically sample the interim readout $\boldsymbol{r}$ in the first stage. Successively, we decode $\boldsymbol{r}$ with some deterministic algorithms in the second stage and attempt to improve it by taking advantage of the information involved in the syndrome that is derived from $\boldsymbol{r}$. If the inferred state $\boldsymbol{z}^{*}$ is better than $\boldsymbol{r}$, we select $\boldsymbol{z}^{*}$ as the estimate of $\boldsymbol{z}$. We can iterate this post-readout decoding until saturation, that is, no further improvement occurs. Therefore, our algorithm can be considered as a hybrid computation of earlier stochastic and later deterministic algorithms, as shown schematically in Fig. 5. If the stochastic algorithm is executed by a quantum computer including a quantum annealer, our algorithm is merely a hybrid quantum and classical computation. We studied how such decoding can be implemented from the classical viewpoint. We analyzed the similarities and differences between decoding for the PE scheme and QAC. For deterministic decoding in the second stage, we focus on one-step MVD, which is the simplest conventional decoding method for classical repetition codes. We compare the one-step MVD for QAC and that for the PE scheme. The one-step MVD relies on a priori probability (APP) decoding, which resembles MPM decoding. Its theoretical foundation was first described by Massey Massey (1963, 1968), who considered a special set of syndromes called orthogonal parity checks and their correlation with errors. Massey’s theory suggests that we should be careful to compare the correct set of readouts of the physical states in order to perform fair majority voting and derive the correct inference. Let us briefly review Massey’s theory, called threshold decoding, using spin representation.
@@
holds for any code state $\boldsymbol{z}$. Thus, $A_{mi}$ is a function of only $\boldsymbol{e}$. $\boldsymbol{A}_{m}$ has another notable property: every $A_{mi}$ involves the noise component $e_{m}$, but no other noise component $e_{i}$ $\left(i\neq m\right)$ is involved in more than one syndrome in the set, i.e., every $e_{i}$ $\left(i\neq m\right)$ appears in only one of the sets $\left\{C_{1}\left(m\right),\ldots,C_{J}\left(m\right)\right\}$. Such a subset $\boldsymbol{A}_{m}$ of composite syndromes is said to be orthogonal on $e_{m}$. Massey showed that a set $\boldsymbol{A}_{m}$ of syndromes orthogonal on $e_{m}$ is a good APP estimator for $e_{m}$ Massey (1963, 1968).

### 28 Feb 2024 v3

@@ I Introduction
In the ME scheme, each fully connected spin (logical spin) is encoded on an associated chain of physical spins, and physical spins belonging to the same chain are connected by strong ferromagnetic interactions to align them energetically. Long-range interactions between logical spins are replaced by short-range interactions between chains of physical spins. As a result, a general Ising problem that is defined on a fully connected graph of many nodes can be described by an equivalent problem that is defined on a sparse graph with a larger number of nodes Vinci et al. (2015). In contrast, in the PE scheme, which was first proposed by Sourlas Sourlas (2005) and later by Lechner, Hauke, and Zoller (LHZ) Lechner et al. (2015), the parity of each pair of logical spins is encoded on an associated physical spin that is arranged on a two-dimensional lattice. Four nearest-neighbor physical spins are coupled by four-body interactions and constitute a PE spin network. Although this approach can avoid long-range interactions, it requires four-body interactions, which appears physically unnatural and unfeasible. However, Puri et al. later suggested that the PE spin network is realizable by mapping it to a network of two-photon-driven Kerr parametric oscillators (KPOs) with only local interactions, where a four-body interaction is mapped to a four-wave mixing of four-mode oscillator fields in the Josephson junction coupler Puri et al. (2017). Such a network of KPOs can be implemented using a circuit QED platform employing superconducting devices, which has been studied extensively to realize quantum computing with the current noisy intermediate-scale quantum devices Albash et al. (2016); Lechner et al. (2015). The KPOs have been experimentally realized by using trapped ions Ding et al. (2017) and Josephson parametric oscillators Grimm et al. (2020); Frattini et al. ; Wang et al. (2019); Yamaji et al. (2022). A network of two KPOs has been recently realized using a static capacitive coupling Yamaji et al. (2023). Chancellor, Zohren, and Warburton independently designed a four-body coupling of a superconducting flux quantum bit (qubit), which is a more naive model of the quantum spin, using additional ancillary qubits Chancellor et al. (2017). In the PE scheme, the optimization problem is encoded in the dataset of the magnitudes as well as orientations of the local fields that act independently on each physical spin, whereas problem-independent four-body couplings among the neighboring physical spins enforce the parity constraints.
@@
In this study, we investigated another aspect of the PE scheme, that is, its error-correcting capability, by comparing it with that of the quantum annealing correction (QAC) scheme Jordan et al. (2006); Young et al. (2013a); Pudenz et al. (2014, 2015); Vinci et al. (2015); Bookatz et al. (2015); Matsuura et al. (2016, 2017); Pearson et al. (2019). QAC is a QA system that is merged with the classical repetition code to mitigate the errors arising in the QA algorithm. Its error-correcting capability has been discussed in terms of that of the classical repetition code. The error-correcting capability of the PE scheme is more subtle. LHZ suggested that the PE scheme includes intrinsic fault tolerance because the logical spins are encoded redundantly and nonlocally in the physical spins. Pastawski and Preskill suggested that the PE scheme can be viewed as a classical low-density parity-check (LDPC) code that makes the scheme highly robust against weakly correlated spin-flip noise Pastawski and Preskill (2016). However, a numerical study using simulated QA and parallel tempering showed no clear evidence that intrinsic fault tolerance can be used to improve the sampling efficiency of the ground state of logical spins Albash et al. (2016). To elucidate this inconsistency, we conducted discrete-time Markov chain Monte Carlo (DT-MCMC) simulations to sample the sequence of state of physical spins under various constant simulation parameters. As our MCMC is reversible, the sequence converges with the stationary distribution related to the Boltzmann distribution. Through the MCMC simulation, we stored the states of the physical spins sampled during the MCMC iteration in the buffer memory. After completing the simulation, we decoded the physical state using one-step majority vote decoding (one-step MVD) Massey (1962, 1968); Laferriere (1977); Clerk and Cain (1981); Shu and Costello (2004) to infer the state of the logical spins and calculated their energies for all the elements in the stored sequence of the physical states. Their energy spectra indicated that a substantial number of non-code states contributed to the logical ground state after the decoding. In particular, we found that the optimal simulation parameters for sampling the physical ground state using no post-readout decoding and the logical ground states using one-step MVD are very different.
@@ II.1 Soft annealing
$\stackrel{{\scriptstyle[}}{{j}}=1]{N}{\bigoplus}H_{ij}c_{j}=0\;\left(i=1,\ldots,N-K\right).$ (2)
@@
This defines the Boolean matrix $\boldsymbol{M}$ called the “parity constraint matrix” that is associated with the parity check matrix $\boldsymbol{H}$ of the code. $M_{k_{1}\cdots k_{q}}^{l}$ is $1$ if each parity equation consists of the Boolean sum of $c_{k_{l}}\left(l=1,\ldots,p\right)$. The physical state $\boldsymbol{z}$ that satisfies Eq. (4) is called a code state. For the general physical state $\boldsymbol{z}^{\prime}$, we refer to $S_{i}\left(\boldsymbol{z}^{\prime}\right)$ syndrome, which is the classical counterpart of the stabilizer operator defined in terms of the quantum correcting codes Gottesman (1996); Jordan et al. (2006). In some studies, it is also called the parity check in the context of the communication application Massey (1962, 1968).
@@
$P\left(\boldsymbol{z}|\boldsymbol{J}\right)$ $=$ $\frac{P\left(\boldsymbol{J}|\boldsymbol{z}\right)P\left(\boldsymbol{z}\right)P\left(\boldsymbol{Z}\right)}{\sum_{\boldsymbol{z}}P\left(\boldsymbol{J}|\boldsymbol{z}\right)P\left(\boldsymbol{z}\right)P\left(\boldsymbol{Z}\right)}$ (5)
@@
where $\kappa$ is a constant independent of $\boldsymbol{z}$, which is to be determined by the normalization condition $\mathop{\sum_{\boldsymbol{z}}P\left(\boldsymbol{z}|\boldsymbol{J}\right)=1}$. $P\left(\boldsymbol{z}\right)$ and $P\left(\boldsymbol{Z}\right)$ are the prior probabilities for the physical spin state $\boldsymbol{z}$ and logical spin state $\boldsymbol{Z}$, respectively Vicente et al. (2003). In the present scenario, we assume an informative prior for the parity constraints of the physical spin state $\boldsymbol{z}$ but ignorance prior for the logical spin state $\boldsymbol{Z}$, that is,
@@ II.4 QAC
which commutes with $\hat{H}^{enc}\left(\hat{\boldsymbol{z}}\right)$ and is introduced to take advantage of the prior knowledge of $K-1$ sets of the parity constraints on the $K$ physical spin variables $\hat{z}_{i\,k}$ for $k=1,\ldots,K$ in the Bayes inference. $\hat{H}^{pen}\left(\hat{\boldsymbol{z}}\right)$ enforces the alignment of the $K$ physical spins in each logical group. The Hamiltonian $\hat{H}^{two}\left(\hat{\boldsymbol{z}}\right)$ introduces correlations between the physical spins in different logical groups, which are absent in the Hamiltonian $\hat{H}^{phys}\left(\hat{\boldsymbol{z}}\right)$ in the communication scenario. In the limit $\gamma\rightarrow 0$, $\hat{H}^{QAC}\left(\hat{\boldsymbol{z}}\right)$ approaches the Hamiltonian for $K$ independent copies of the $N$ spin system, each of which is described by the logical Hamiltonian $\left.\hat{H}^{logi}\left(\hat{\boldsymbol{Z}}\right)\right|_{\hat{Z}_{i}\mapsto\hat{z}_{i,k}}$. Hence, we refer to a set of $N$ physical spins specified by $k$ as a replica. In contrast, in the limit $\beta\rightarrow 0$, $\hat{H}^{QAC}\left(\hat{\boldsymbol{z}}\right)$ approaches the Hamiltonian $\hat{H}^{pen}\left(\hat{\boldsymbol{z}}\right)$ that energetically penalizes the non-code state of the $\left[N,1\right]$ repetition spin-flip code. In this case, it should be noted that if we select the weight-two stabilizer in Eq. (38), $\hat{H}^{QAC}\left(\hat{\boldsymbol{z}}\right)$ is formally identical to that in simulated QA (SQA) using the path-integral Monte Carlo method Martoňák et al. (2002); Heim et al. (2015); Waidyasooriya and Hariyama (2020); Hu and Wang (2021). This implies that SQA may be re-investigated from the viewpoint of the classical error-correcting code. We can describe the quantum dynamics of QAC by combining the standard transverse-field driver Hamiltonian based on local Pauli-$x$ operators
@@ III Decoding of post-readout spin state
$\boldsymbol{Z}^{*}=\underset{\boldsymbol{Z}\in\left\{\pm 1\right\}^{K}}{\arg\max}P\left(\boldsymbol{Z}|\boldsymbol{J}\right)=\underset{\boldsymbol{Z}\in\left\{\pm 1\right\}^{K}}{\arg\max}L\left(\boldsymbol{Z},\boldsymbol{J}\right)=\underset{\boldsymbol{Z}\in\left\{\pm 1\right\}^{K}}{\arg\min}H^{logi}\left(\boldsymbol{Z}\right)=\boldsymbol{Z}_{g}$ (47)
@@
$\boldsymbol{z}^{*}=\underset{\boldsymbol{z}\in\left\{\pm 1\right\}^{N}}{\arg\max}P\left(\boldsymbol{z}|\boldsymbol{J}\right)=\underset{\boldsymbol{z}\in\left\{\pm 1\right\}^{N}}{\arg\max}L\left(\boldsymbol{z},\boldsymbol{J}\right)=\underset{\boldsymbol{z}\in\left\{\pm 1\right\}^{N}}{\arg\min}H^{phys}\left(\boldsymbol{z}\right)=\boldsymbol{z}_{g}$ (48)
@@
In the following, we focus on the case where the target states are the logical ground states $\boldsymbol{Z}_{g}$ and associated physical code state $\boldsymbol{z}_{g}=\boldsymbol{Z}_{g}\otimes\boldsymbol{Z}_{g}$. This is because our main objective is the development of an efficient solver for combinatorial optimization problems by taking advantage of nature-inspired approaches. It is assumed that we have a method that can statistically sample the neighboring states of the physical target state $\boldsymbol{r}\sim\boldsymbol{z}_{g}$, such as simulated annealing based on Monte Carlo simulation and QA. The readout state $\boldsymbol{r}$ may be a non-code state. We address the decoding of the interim readout $\boldsymbol{r}$ to infer the target state $\boldsymbol{z}^{*}$ by taking advantage of the informative prior for parity constraints incorporated by the error-correcting codes. Let us consider the following scenario. Given $\boldsymbol{J}$ and $\boldsymbol{h}$, we statistically sample the interim readout $\boldsymbol{r}$ in the first stage. Successively, we decode $\boldsymbol{r}$ with some deterministic algorithms in the second stage and attempt to improve it by taking advantage of the information involved in the syndrome that is derived from $\boldsymbol{r}$. If the inferred state $\boldsymbol{z}^{*}$ is better than $\boldsymbol{r}$, we select $\boldsymbol{z}^{*}$ as the estimate of $\boldsymbol{z}$. We can iterate this post-readout decoding until saturation, that is, no further improvement occurs. Therefore, our algorithm can be considered as a hybrid computation of earlier stochastic and later deterministic algorithms, as shown schematically in Fig. 5. If the stochastic algorithm is executed by a quantum computer including a quantum annealer, our algorithm is merely a hybrid quantum and classical computation. We studied how such decoding can be implemented from the classical viewpoint. We analyzed the similarities and differences between decoding for the PE scheme and QAC. For deterministic decoding in the second stage, we focus on one-step MVD, which is the simplest conventional decoding method for classical repetition codes. We compare the one-step MVD for QAC and that for the PE scheme. The one-step MVD relies on a priori probability (APP) decoding, which resembles MPM decoding. Its theoretical foundation was first described by Massey Massey (1962, 1968), who considered a special set of syndromes called orthogonal parity checks and their correlation with errors. Massey’s theory suggests that we should be careful to compare the correct set of readouts of the physical states in order to perform fair majority voting and derive the correct inference. Let us briefly review Massey’s theory, called threshold decoding, using spin representation.
@@
holds for any code state $\boldsymbol{z}$. Thus, $A_{mi}$ is a function of only $\boldsymbol{e}$. $\boldsymbol{A}_{m}$ has another notable property: every $A_{mi}$ involves the noise component $e_{m}$, but no other noise component $e_{i}$ $\left(i\neq m\right)$ is involved in more than one syndrome in the set, i.e., every $e_{i}$ $\left(i\neq m\right)$ appears in only one of the sets $\left\{C_{1}\left(m\right),\ldots,C_{J}\left(m\right)\right\}$. Such a subset $\boldsymbol{A}_{m}$ of composite syndromes is said to be orthogonal on $e_{m}$. Massey showed that a set $\boldsymbol{A}_{m}$ of syndromes orthogonal on $e_{m}$ is a good APP estimator for $e_{m}$ Massey (1962, 1968).